OpenStudy (anonymous):

Find the quadratic equation for these given solutions: x = 2i and x = -2i

OpenStudy (anonymous):

(x-2i)(x+2i)=0 x^2+4=0

OpenStudy (a_clan):

x^2 - Sx + P = 0 where S= a+b and P=a.b where a, b are the roots of quadratic equation.

OpenStudy (anonymous):

(x-a)*(x-b)=0 where a and b are roots